Hello-
I'm stuck on an Algebra II problem and hope you can help. The equation is as follows:
1/8 (3m-5)= 6/5 (6m-2)+5
I have never had a problem like this one.
I would appreciate your help and guidance.
Thank you very much.
Sam

Answer by stanbon(75887)

1/8 (3m-5)= 6/5 (6m-2)+5
Multiply both sides by 40 to get:
5(3m-5) = 48(6m-2) + 200
15m-25 = 288m-96 + 200
-273m=25-96+200
-273m=129
m=-0.4725...
Cheers,
Stan H.